# SReferences

### 25 Mar 2024 v1

@@ Abstract
In periodically driven (Floquet) systems, evolution typically results in an infinite-temperature thermal state due to continuous energy absorption over time. However, before reaching thermal equilibrium, such systems may transiently pass through a meta-stable state known as a prethermal state. This prethermal state can exhibit phenomena not commonly observed in equilibrium, such as discrete time crystals (DTCs), making it an intriguing platform for exploring out-of-equilibrium dynamics. Here, we investigate the relaxation dynamics of initially prepared product states under periodic driving in a kicked Ising model using the IBM Quantum Heron processor, comprising 133 superconducting qubits arranged on a heavy-hexagonal lattice, over up to $100$ time steps. We identify the presence of a prethermal regime characterised by magnetisation measurements oscillating at twice the period of the Floquet cycle and demonstrate its robustness against perturbations to the transverse field. Our results provide evidence supporting the realisation of a period-doubling DTC in a two-dimensional system. Moreover, we discover that the longitudinal field induces additional amplitude modulations in the magnetisation with a period incommensurate with the driving period, leading to the emergence of discrete time quasicrystals (DTQCs). These observations are further validated through comparison with tensor-network and state-vector simulations. Our findings not only enhance our understanding of clean DTCs in two dimensions but also highlight the utility of digital quantum computers for simulating the dynamics of quantum many-body systems, addressing challenges faced by state-of-the-art classical simulations.
@@
Periodically driven (Floquet) systems host novel phases of matter inaccessible in thermal equilibrium. Notably, discrete time crystals (DTCs) [1, 2, 3, 4, 5] represent genuine out-of-equilibrium phases of matter [6, 7, 8, 9] feasible in Floquet systems [10, 11, 12]. A DTC is characterised by subharmonic responses breaking discrete time-translational symmetry imposed by the periodic drive. However, sustaining DTCs as transient meta-stable states faces challenges due to thermalisation, where many-body interactions drive low-entangled states to highly entangled, high-energy states. Overcoming this obstacle requires imparting a many-body localised nature to the dynamics.
@@
One strategy to circumvent rapid thermalisation in driven systems is by introducing disorder in the Floquet Hamiltonian, inducing many-body localisation (MBL) to break ergodicity [11, 12, 13, 14, 15]. Recently, disorder-induced MBL-based DTCs (MBL-DTCs) have been demonstrated on digital quantum computers in one dimension [16, 17, 18]. Furthermore, topological time crystalline order has been achieved in a periodically driven disordered toric code on a superconducting quantum computer [19]. Another avenue for DTCs involves the prethermal regime of periodically driven clean systems in two or higher dimensions [20, 21, 22, 23, 24, 25, 26, 27]. Unlike MBL-DTCs, prethermal DTCs are not stabilised in one-dimensional systems with short-range interactions, aligning with the absence of symmetry breaking at finite temperatures in one dimension. Therefore, realising a clean prethermal DTC requires two or higher dimensions, or otherwise long-range interactions.
@@
In simulating dynamics of quantum many-body systems in two dimensions, tensor-network methods have been extensively utilised for large systems beyond the capabilities of state-vector simulations [28, 29, 30, 3]. However, accurate tensor-network simulations over extended periods become challenging in two dimensions due to breakdowns in low-rank tensor approximations when entanglement exceeds certain thresholds dictated by bond dimensions. Conversely, recent advancements in noisy intermediate-scale quantum devices have introduced digital quantum computers as another tool to investigate out-of-equilibrium phases of matter, including DTCs. Indeed, a recent study has observed indications of clean DTCs in two dimensions using a digital quantum computer [32]. Nevertheless, the previous study has been performed on systems of a few tens of qubits, leaving large-scale digital quantum simulations, comparable to state-of-the-art classical tensor-network simulations, of clean DTCs to examine their stability in two dimensions.
@@
Here, we demonstrate the realisation of clean DTCs on a two-dimensional heavy-hexagonal lattice of $133$ qubits (see Fig. 1a) using an IBM Quantum Heron processor, ibm_torino. By applying periodic driving to initial product states in a kicked Ising model, involving both transverse and longitudinal fields, we measure local magnetisation to observe its subharmonic response. With a simple error mitigation protocol based on a depolarising noise model, our results are first validated by showing agreement with both tensor-network simulations of the $133$-qubit system and state-vector simulations of a $28$-qubit system for up to $50$ time steps. We then observe a subharmonic period-doubling response of local magnetisation persisting for at least $100$ time steps, confirming its stability against perturbations to the transverse field, which thereby provides evidence for a realisation of clean DTCs in two dimensions. Furthermore, we observe other longer-period subharmonic responses with frequencies incommensurate with the driving period, thus identified as discrete time quasicrystals (DTQCs) [22, 25].
@@
$\hat{H}(t)=\left\{\,\begin{aligned} &h_{x}\sum_{i=0}^{L-1}\hat{X}_{i},&\text{ for }0\leqslant t<T/2\\ &h_{z}\sum_{i=0}^{L-1}\hat{Z}_{i}-J\sum_{\langle i,j\rangle}\hat{Z}_{i}\hat{Z}_{j}&\text{ for }T/2\leqslant t<T\end{aligned}\right.,$ (1)
@@
where $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})=\exp\left[-\mathrm{i}\theta_{J}\hat{Z}_{i}\hat{Z}_{j}/2\right]$, $\hat{R}_{Z_{i}}(\theta_{z})=\exp\left[-\mathrm{i}\theta_{z}\hat{Z}_{i}/2\right]$, and $\hat{R}_{X_{i}}(\theta_{x})=\exp\left[-\mathrm{i}\theta_{x}\hat{X}_{i}/2\right]$ are $ZZ$, $Z$, and $X$ rotation gates with rotation angles $\theta_{J}=-JT$, $\theta_{z}=h_{z}T$, and $\theta_{x}=h_{x}T$, respectively. Since each qubit is coupled to at most three adjacent qubits on the heavy-hexagonal lattice, operation of all the two-qubit gates $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ has to be divided into three layers (see Fig. 1b). Each layer consists of $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ gates on red, blue, or green edges in Fig. 1a, allowing for parallel operation.
@@
The time-evolved state at stroboscopic times $t=nT$ with integer $n$ is expressed as $|\psi(t)\rangle=(\hat{U}_{\rm F})^{n}|\psi(0)\rangle,$ where $|\psi(0)\rangle$ represents the initial state. Our primary focus lies in measuring local magnetisation defined as $\langle\hat{Z}_{j}(t)\rangle=\langle\psi(t)|\hat{Z}_{j}|\psi(t)\rangle,$ where $\hat{Z}_{j}(t)=(\hat{U}_{\rm F}^{\dagger})^{n}\hat{Z}_{j}(\hat{U}_{\rm F})^{n}$ is the Heisenberg representation of the Pauli $\hat{Z}_{j}$ operator, and $\langle\cdots\rangle=\langle\psi(0)|\cdots|\psi(0)\rangle$ denotes the expectation value with respect to the initial state. The initial state is prepared as a product state in the computational basis, forming a stripe pattern of $|0\rangle$’s and $|1\rangle$’s, represented by white and black circles in Fig. 1a. Among the three independent model parameters, we set $\theta_{J}=-\pi/2$ and vary the other two parameters $\theta_{x}$ and $\theta_{z}$. The gate $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ at $\theta_{J}=-\pi/2$ is decomposed into the CZ gate, the native two-qubit gate of ibm_torino, and the $S$ gate, as $\hat{R}_{Z_{i}Z_{j}}(-\pi/2)=e^{i\pi/4}\widehat{\rm CZ}_{ij}(\hat{S}_{i}^{\dagger}\otimes\hat{S}_{j}^{\dagger})$.
@@
Figure 1: Two-qubit gate connectivity on a heavy-hexagonal lattice and initial product states. a, The overall device geometry of ibm_torino, comprising a hevey-hexagonal lattice of $L=133$ qubits. Each circle represents a qubit and the edges indicate the qubit connectivity. Three layers of $R_{ZZ}$ gates in a single Floquet circle are highlighted in red, blue, and green (also see b). The enclosed area marked by the yellow line represents the system of $L=28$ qubits, utilised for comparison with state-vector simulations. The initial state is prepared as a product state with qubits denoted by white (black) circles initialised to be $|0\rangle$ ($|1\rangle$). b, Schematic representation of the single-cycle Floquet operator $\hat{U}_{\rm F}$. The red, blue, and green boxes correspond to the three layers of $R_{ZZ}$ gates, each applied in parallel, indicated in a, while the white boxes represent the products of $R_{X}$ and $R_{Z}$ gates. The horizontal lines represent the qubits on which the quantum gates operate.
@@
For convenience, we introduce a perturbation parameter $\epsilon$ to the transverse field as $2\epsilon=\pi-\theta_{x}.$ When $\varepsilon=0$, the dynamics of $\hat{Z}_{j}(t)$ becomes trivial because a single Floquet cycle simply flips the sign of the local magnetisation, $\hat{Z}_{j}(t+T)=\hat{U}_{\rm F}\hat{Z}_{j}(t)\hat{U}_{\rm F}^{\dagger}=(-1)\hat{Z}_{j}(t)$. This demonstrates that a period-doubling DTC with $|\langle\hat{Z}_{j}(t)\rangle|=1$ is realised, at least at the fine-tuned parameter $\theta_{x}=\pi$. Our primary interest therefore lies in the subharmonic response of the magnetisation for $\epsilon>0$.
We utilise ibm_torino, the IBM Quantum Heron processor comprising $133$ superconducting qubits arranged on the heavy-hexagonal lattice (see Fig. 1a) [33]. The median infidelity of the native two-qubit gates (i.e., CZ gates) is approximately $4\times 10^{-3}$, while the infidelity of single-qubit gates is around $4\times 10^{-4}$ with the median read-out error of $1.7\times 10^{-2}$ (see also Supplementary Information S1). Each Floquet cycle involves $150$ CZ gates for the $L=133$ system, totaling $150n$ CZ gates for $n$ time steps. Given the three non-parallelisable layers of two-qubit gates per cycle, the circuit depth for the $n$ time steps is $3n$. Considering a two-qubit gate time of about $100$ns, the real-time duration from state preparation to final measurement for the maximum circuit depth $300$ for $n=100$ is estimated as roughly $30\mu$s, significantly shorter than the median single-qubit coherence times $T_{1}=165\mu$s and $T_{2}=138\mu$s.
@@
First, we introduce an error-mitigation scheme to validate that the quantum device provides reliable results for magnetisation dynamics. We measure the time evolution of the averaged magnetisation over a set of qubits,
@@
where $|A|$ denotes the number of qubits in set $A$, and we choose that $A=\{63,64,65,66,67,68,69,70,71\}$ for the $L=28$ system and $A=\{57,58,59,60,61,62,63,64,65,66,67,68,69,70,71\}$ for the $L=133$ system (see Fig. 1a). In utilising the quantum device, we estimate the expectation value of ${\hat{Z}_{\rm avg}}(t)$ at each time step by computing the sample mean of outcomes from projective measurements on all qubits within $A$ in the computational basis over $2^{14}$ samples. The statistical error associated with this estimate is determined as the sample standard deviation of the mean. The results are shown in Fig. 2a,c, where neither error-suppression methods such as dynamical decoupling [34, 35] nor error-mitigation methods such as zero-noise extrapolation [36, 37] and probabilistic error cancellation [38] are used (the same holds for the other results presented below).
Figure 2: Error-mitigation protocol and comparison with classical simulations. a, Raw data of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ (cyan circles) and $f(\theta_{x}=\pi)=|\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}|$ at $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$ (black crosses) for $L=28$. b, Error-mitigated data (yellow diamonds) of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ for $L=28$. Numerically exact results obtained by state-vector simulations are also shown with black squares in b. c-d, Same as a-b but for $L=133$. Results of classical tensor-network simulations with bond dimensions $\chi=100$ and 200 are also shown with green and blue triangles, respectively, in d. Error bars in b and d represent the propagated error due to sampling errors in the quantities in the numerator and the denominator of Eq. (4).
As described above, at $\theta_{x}=\pi$, the noiseless expectation value satisfies $|\langle\hat{Z}_{\rm avg}(t)\rangle|=1$. However, the absolute values of the raw data obtained from the quantum device, denoted as $f(\theta_{x}=\pi):=|\langle\hat{Z}_{\rm avg}(t)\rangle_{0}|$, deviate from the ideal value $1$, with the deviation increasing over time steps, as observed in Fig. 2a,c. To account for this signal decay, we introduce a global depolarising noise model, where the expectation value $\langle\hat{O}(t)\rangle_{0}$ of an observable $\hat{O}$ subject to depolarising noise is given by $\langle\hat{O}(t)\rangle_{0}=f\langle\hat{O}(t)\rangle+(1-f){\rm Tr}[\hat{O}(t)]/2^{L}$ [39, 40]. Here, $f$ is a parameter that characterises the depolarising noise model, with $\langle\hat{O}(t)\rangle$ representing the ideal expectation value of $\hat{O}$ and ${\rm Tr}[\hat{O}(t)]/2^{L}$ being the expectation value over the maximally mixed state. Generally, $f$ depends on both the circuit and observable, i.e., $f=f(\theta_{J},\theta_{x},\theta_{z},n,\hat{O})$. Since $|\langle\hat{Z}_{\rm avg}(t)\rangle|=1$ at $\theta_{x}=\pi$ and ${\rm Tr}[\hat{Z}_{\rm avg}(t)]=0$ as $\hat{Z}_{j}(t)$ is traceless, the parameter $f$ can be estimated in this trivial case as $f(\theta_{J},\pi,\theta_{z},n,\hat{Z}_{\rm avg})=|\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}|$, where $\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}$ is $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ obtained at $\theta_{x}=\pi$. For general $\theta_{x}$, it is difficult to estimate $f$ because the ideal expectation value $\langle\hat{Z}_{\rm avg}(t)\rangle$ is not available. To circumvent this issue, we approximate $f(\theta_{J},\theta_{x},\theta_{z},n,\hat{Z}_{\rm avg})$ by $f(\theta_{J},\pi,\theta_{z},n,\hat{Z}_{\rm avg})=f(\theta_{x}=\pi)$. This approximation leads to an error mitigation scheme:
@@
Similar error-mitigation protocols have been successfully applied previously to correct magnetisation [18] and out-of-time-ordered correlators [41].
@@
The same error mitigation scheme demonstrates excellent performance even for the $L=133$ system, as shown in Fig. 2c,d. In parallel, we employ a two-dimensional tensor-network state (2dTNS) method as a classical counterpart (Supplementary Information S5). The 2dTNS results presented here converge with respect to the bond dimension $\chi$, which governs the accuracy of the approximation inherent in the 2dTNS method, for time steps up to at least 50 (see Supplementary Information S3 for further comparisons with longer time steps and different parameters). Once again, the remarkable agreement between the error-mitigate data and the converged tensor-network simulation results confirms the reliability of the quantum device outcomes. This validation firmly establishes digital quantum simulations on the quantum device as a compelling tool to explore clean DTCs in two dimensions. It should also be emphasised here that the number of time steps achievable in this quantum device, providing reliable results along with the simple mitigation protocol, easily exceeds that of previous similar dynamics experiments using the IBM Eagle processor of 127 qubits with the median infidelity of two-qubit gates $\sim 1\times 10^{-2}$ [42], where at most 20 time steps were evolved, albeit with various error-mitigation techniques involved.
@@
Having validated the reliability of quantum hardware results, we now delve into discrete time-crystalline orders in two dimensions. Figure 3 shows the long-time dynamics, spanning up to 100 time steps, of the raw and mitigated magnetisation, $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ and $\langle\hat{Z}_{\rm avg}(t)\rangle$, respectively, on the heavy-hexagonal lattice of $L=133$ qubits for various sets of parameters $(\theta_{x},\theta_{z})$. Overall, the decay of the magnetisation becomes more pronounced as $\theta_{x}$ decreases. Specifically, period-doubling oscillations persist even around $t/T=100$ for $\theta_{x}\geqslant 0.8$, while they are barely observable for $t/T\gtrsim 20$ at $\theta_{x}=0.7$, suggesting thermalisation. These observations lead to the conclusion that DTCs observed on the heavy-hexagonal lattice remain stable in the range $0.8\pi\leqslant\theta_{x}\leqslant\pi$, where a prethermal plateau [23] is distinctly visible withing the time steps $0\leqslant t/T\leqslant 100$. This is in sharp contrast to the behavior in one dimension, where magnetisation oscillations quickly decay, irrespective of the parameters $(\theta_{x},\theta_{z})$ away from the trivial point at $\theta_{x}=\pi$, as we have also confirmed in Supplementary Information S4 using the same quantum device.

### 12 Mar 2026 v2

# Unveiling clean two-dimensional discrete time crystals on a digital quantum computer

@@ Abstract
In periodically driven (Floquet) systems, evolution typically results in an infinite-temperature thermal state due to continuous energy absorption over time. However, before reaching thermal equilibrium, such systems may transiently pass through a meta-stable state known as a prethermal state. This prethermal state can exhibit phenomena not commonly observed in equilibrium, such as discrete time crystals (DTCs), making it an intriguing platform for exploring out-of-equilibrium dynamics. Here, we investigate the relaxation dynamics of initially prepared product states under periodic driving in a kicked Ising model using the IBM Quantum Heron processor, comprising 133 superconducting qubits arranged on a heavy-hexagonal lattice, over up to $100$ time steps. We identify a clean two-dimensional DTC characterised by magnetisation measurements oscillating at twice the period of the Floquet cycle and demonstrate its robustness against perturbations to the transverse field. This stability does not rely on many-body localisation or on high-frequency Floquet prethermalisation, but emerges in a clean, disorder-free setting. Moreover, we discover that the longitudinal field induces additional amplitude modulations in the magnetisation with a period incommensurate with the driving period, leading to the emergence of an incommensurately modulated discrete time-crystal (IM-DTC) response. These observations are further validated through comparison with tensor-network and state-vector simulations. Our findings not only provide insight into clean DTC and IM-DTC dynamics in two dimensions but also highlight the utility of gate-based quantum computers for simulating the dynamics of quantum many-body systems, complementing state-of-the-art classical simulations in regimes where entanglement growth challenges their convergence.
@@
Periodically driven (Floquet) systems host novel phases of matter inaccessible in thermal equilibrium. Notably, discrete time crystals (DTCs) [1, 2, 3, 4, 5, 6] represent genuine out-of-equilibrium phases of matter [7, 8, 9] feasible in Floquet systems [10, 11, 12]. A DTC is characterised by subharmonic responses breaking discrete time-translational symmetry imposed by the periodic drive. However, sustaining DTCs as transient meta-stable states faces challenges due to thermalisation, where many-body interactions drive low-entangled states to highly entangled, high-energy states. Overcoming this obstacle requires imparting a many-body localised nature to the dynamics.
@@
One strategy to circumvent rapid thermalisation in driven systems is by introducing disorder in the Floquet Hamiltonian, inducing many-body localisation (MBL) to break ergodicity [11, 12, 13, 14, 15]. Recently, disorder-induced MBL-based DTCs (MBL-DTCs) have been demonstrated on digital quantum computers in one dimension [16, 17, 18, 19]. Furthermore, topological time crystalline order has been achieved in a periodically driven disordered toric code on a superconducting quantum computer [20, 21]. Another avenue for DTCs involves the prethermal regime of periodically driven clean systems in two or higher dimensions [22, 23, 24, 25, 26, 27, 28, 29, 30]. Unlike MBL-DTCs, prethermal DTCs are not stabilised in one-dimensional systems with short-range interactions, aligning with the absence of symmetry breaking at finite temperatures in one dimension. Therefore, realising a clean DTC requires two or higher dimensions, or otherwise long-range interactions.
Another avenue for DTCs involves periodically driven clean systems operated in a Floquet prethermal regime, where a sufficiently large driving frequency gives rise to long-lived time-crystalline responses governed by an effective static Hamiltonian [22, 25, 26, 30]. In addition, it has been demonstrated that robust DTC behaviour can also occur in clean, strongly interacting systems at finite driving frequencies, beyond the strict high-frequency limit and without relying on MBL protection [23, 27, 28, 29]. In such disorder-free settings, one-dimensional systems tend to thermalise, whereas higher-dimensional lattices or effectively long-range couplings can sustain substantially longer-lived time-crystalline behaviour. This suggests that realising a clean DTC generically requires either two or higher spatial dimensions, or sufficiently long-range interactions, and our results below provide an explicit realisation of the former scenario on a two-dimensional heavy-hexagonal lattice at finite driving frequency, without invoking MBL or high-frequency Floquet prethermalisation.
@@
In simulating dynamics of quantum many-body systems in two dimensions, tensor-network methods have been extensively utilised for large systems beyond the capabilities of state-vector simulations [31, 32, 33, 34, 35]. However, accurate tensor-network simulations over extended periods become challenging in two dimensions due to breakdowns in low-rank tensor approximations when entanglement exceeds certain thresholds dictated by bond dimensions. Conversely, recent advancements in noisy intermediate-scale quantum devices have introduced digital quantum computers as another tool to investigate out-of-equilibrium phases of matter, including DTCs. Despite these advances, digital quantum simulations of DTCs in genuinely two-dimensional geometries with system sizes comparable to state-of-the-art classical tensor-network simulations remain largely unexplored, and a systematic examination of their stability in such large-scale settings is still lacking.
@@
Here, we demonstrate the realisation of clean DTCs on a two-dimensional heavy-hexagonal lattice of $133$ qubits (see Fig. 1a) using an IBM Quantum Heron processor, ibm_torino. By applying periodic driving to initial product states in a kicked Ising model [36], involving both transverse and longitudinal fields, we measure local magnetisation to observe its subharmonic response. With a simple error mitigation protocol based on a depolarising noise model, our results are first validated by showing agreement with both tensor-network simulations of the $133$-qubit system and state-vector simulations of a $28$-qubit system for up to $50$ time steps. We then observe a subharmonic period-doubling response of local magnetisation persisting for at least $100$ time steps, confirming its stability against perturbations to the transverse field, which thereby provides evidence for a realisation of prethermal DTCs in two dimensions. Furthermore, we observe other longer-period subharmonic responses with frequencies incommensurate with the driving period, which we identify as incommensurately modulated DTC (IM-DTC) responses [24, 27].
@@
$\hat{H}(t)=\left\{\,\begin{aligned} &h_{x}\sum_{i=0}^{L-1}\hat{X}_{i},&\text{ for }0\leq t<T/2\\ &h_{z}\sum_{i=0}^{L-1}\hat{Z}_{i}-J\sum_{\langle i,j\rangle}\hat{Z}_{i}\hat{Z}_{j}&\text{ for }T/2\leq t<T\end{aligned}\right$., (1)
@@
where $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})=\exp\left[-\mathrm{i}\theta_{J}\hat{Z}_{i}\hat{Z}_{j}/2\right]$, $\hat{R}_{Z_{i}}(\theta_{z})=\exp\left[-\mathrm{i}\theta_{z}\hat{Z}_{i}/2\right]$, and $\hat{R}_{X_{i}}(\theta_{x})=\exp\left[-\mathrm{i}\theta_{x}\hat{X}_{i}/2\right]$ are $ZZ$, $Z$, and $X$ rotation gates with rotation angles $\theta_{J}=-JT$, $\theta_{z}=h_{z}T$, and $\theta_{x}=h_{x}T$, respectively. Since each qubit is coupled to at most three adjacent qubits on the heavy-hexagonal lattice, operation of all the two-qubit gates $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ has to be divided into three layers (see Fig. 1b). Each layer consists of $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ gates on red, blue, or green edges in Fig. 1a, allowing for parallel operation. In our implementation on ibm_torino, the Floquet time evolution over $n$ periods is realised as a gate-based (digital) quantum circuit obtained by repeating this sequence of single- and two-qubit gates $n$ times.
@@
The time-evolved state at stroboscopic times $t=nT$ with integer $n$ is expressed as $|\psi(t)\rangle=(\hat{U}_{\mathrm{F}})^{n}|\psi(0)\rangle,$ where $|\psi(0)\rangle$ represents the initial state. Our primary focus lies in measuring local magnetisation defined as $\langle\hat{Z}_{j}(t)\rangle=\langle\psi(t)|\hat{Z}_{j}|\psi(t)\rangle,$ where $\hat{Z}_{j}(t)=(\hat{U}_{\mathrm{F}}^{\dagger})^{n}\hat{Z}_{j}(\hat{U}_{\mathrm{F}})^{n}$ is the Heisenberg representation of the Pauli $\hat{Z}_{j}$ operator, and $\langle\cdots\rangle=\langle\psi(0)|\cdots|\psi(0)\rangle$ denotes the expectation value with respect to the initial state. The initial state is prepared as a product state in the computational basis, forming a stripe pattern of $|0\rangle$’s and $|1\rangle$’s, represented by white and black circles in Fig. 1a. Among the three independent model parameters, we set $\theta_{J}=-\pi/2$ and vary the other two parameters $\theta_{x}$ and $\theta_{z}$. The gate $\hat{R}_{Z_{i}Z_{j}}(\theta_{J})$ at $\theta_{J}=-\pi/2$ is decomposed into the CZ gate, the native two-qubit gate of ibm_torino, and the $S$ gate, as $\hat{R}_{Z_{i}Z_{j}}(-\pi/2)=e^{i\pi/4}\widehat{\mathrm{CZ}}_{ij}(\hat{S}_{i}^{\dagger}\otimes\hat{S}_{j}^{\dagger})$.
@@
Figure 1: Two-qubit gate connectivity on a heavy-hexagonal lattice and initial product states. a, The overall device geometry of ibm_torino, comprising a hevey-hexagonal lattice of $L=133$ qubits. Each circle represents a qubit and the edges indicate the qubit connectivity. Three layers of $R_{ZZ}$ gates in a single Floquet circle are highlighted in red, blue, and green (also see b). The enclosed area marked by the yellow line represents the system of $L=28$ qubits, utilised for comparison with state-vector simulations. The initial state is prepared as a product state with qubits denoted by white (black) circles initialised to be $|0\rangle$ ($|1\rangle$). b, Schematic representation of the single-cycle Floquet operator $\hat{U}_{\mathrm{F}}$. The red, blue, and green boxes correspond to the three layers of $R_{ZZ}$ gates, each applied in parallel, indicated in a, while the white boxes represent the products of $R_{X}$ and $R_{Z}$ gates. The horizontal lines represent the qubits on which the quantum gates operate.
@@
For convenience, we introduce a perturbation parameter $\epsilon$ to the transverse field as $2\epsilon=\pi-\theta_{x}.$ When $\epsilon=0$, the dynamics of $\hat{Z}_{j}(t)$ becomes trivial because a single Floquet cycle simply flips the sign of the local magnetisation, $\hat{Z}_{j}(t+T)=\hat{U}_{\mathrm{F}}^{\dagger}\hat{Z}_{j}(t)\hat{U}_{\mathrm{F}}=(-1)\hat{Z}_{j}(t)$. This demonstrates that a period-doubling DTC with $|\langle\hat{Z}_{j}(t)\rangle|=1$ is realised, at least at the fine-tuned parameter $\theta_{x}=\pi$. Our primary interest therefore lies in the subharmonic response of the magnetisation for $\epsilon>0$.
[…]
@@
First, we introduce a simple baseline error-mitigation scheme, based on a global depolarising-noise model, to validate that the quantum device provides reliable results for magnetisation dynamics. We measure the time evolution of the averaged magnetisation over a set of qubits,
@@
where $|A|$ denotes the number of qubits in set $A$, and we choose that $A=\{63,64,65,66,67,68,69,70,71\}$ for the $L=28$ system and $A=\{57,58,59,60,61,62,63,64,65,66,67,68,69,70,71\}$ for the $L=133$ system (see Fig. 1a). In utilising the quantum device, we estimate the expectation value of ${\hat{Z}_{\rm avg}}(t)$ at each time step by computing the sample mean of outcomes from projective measurements on all qubits within $A$ in the computational basis over $2^{14}$ samples. The statistical error associated with this estimate is determined as the sample standard deviation of the mean. The results are shown in Figs. 2a and 2c, where neither error-suppression methods such as dynamical decoupling [38, 39] nor error-mitigation methods such as zero-noise extrapolation [40, 41] and probabilistic error cancellation [42] are used (the same holds for the other results presented below).
Figure 2: Error-mitigation protocol and comparison with classical simulations. a, Raw data of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ (cyan circles) and $f(\theta_{x}=\pi)=|\langle\hat{Z}_{\rm avg}(t)\rangle_{0,\theta_{x}=\pi}|$ at $(\theta_{x},\theta_{z})=(\pi,0.5\pi)$ (black crosses) for $L=28$. b, Error-mitigated data (yellow diamonds) of the averaged magnetisation $\langle\hat{Z}_{\rm avg}(t)\rangle$ at $(\theta_{x},\theta_{z})=(0.8\pi,0.5\pi)$ for $L=28$. Numerically exact results obtained by state-vector simulations are also shown with black squares in b. c-d, Same as a-b but for $L=133$. Results of classical 2dTNS simulations with bond dimensions $\chi=100$ and 200 are also shown with green and blue triangles, respectively, in d. Error bars in b and d represent the propagated error due to sampling errors in the quantities in the numerator and the denominator of Eq. (4).
[…]
@@
We give more detail for this error-mitigation protocol in Supplementary Information S6. This procedure effectively compensates a global, configuration-independent decay of the signal and does not attempt full noise cancellation. Similar approaches have been successfully applied previously to correct magnetisation [18] and out-of-time-ordered correlators [45].
@@
The same error mitigation scheme demonstrates excellent performance even for the $L=133$ system, as shown in Fig. 2c and 2d. In parallel, we employ a two-dimensional tensor-network state (2dTNS) method as a classical counterpart (Supplementary Information S9). The 2dTNS results presented here converge with respect to the bond dimension $\chi$, which governs the accuracy of the approximation inherent in the 2dTNS method, for time steps up to at least 50 (see Supplementary Information S3 for further comparisons with longer time steps and different parameters). Once again, the remarkable agreement between the error-mitigate data and the converged tensor-network simulation results confirms the reliability of the quantum device outcomes in the regime where the classical simulations remain well controlled. In this sense, gate-based quantum simulations on the device provide a complementary platform to tensor-network methods for exploring DTC dynamics in two dimensions, particularly at intermediate and long times. We additionally remark that the present Heron device allows reliable access to substantially more Floquet steps than reported in earlier experiments on comparable systems, such as studies using the IBM Eagle processor of 127 qubits [46], where evolution was typically limited to about 20 time steps with multiple error-mitigation methods.
@@
Having validated the reliability of quantum hardware results, we now delve into discrete time-crystalline orders in two dimensions. Figure 3 shows the long-time dynamics, spanning up to 100 time steps, of the raw and mitigated magnetisation, $\langle\hat{Z}_{\rm avg}(t)\rangle_{0}$ and $\langle\hat{Z}_{\rm avg}(t)\rangle$, respectively, on the heavy-hexagonal lattice of $L=133$ qubits for various sets of parameters $(\theta_{x},\theta_{z})$. Overall, the decay of the magnetisation becomes more pronounced as $\theta_{x}$ decreases. Specifically, period-doubling oscillations persist even around $t/T=100$ for $\theta_{x}\geqslant 0.8$, while they are barely observable for $t/T\gtrsim 20$ at $\theta_{x}=0.7$, suggesting thermalisation. We observe similar behavior for the fully-polarised initial state of $|0\rangle$’s (see Supplementary information S3). These observations lead to the conclusion that DTCs observed on the heavy-hexagonal lattice remain stable in the range $0.8\pi\leq\theta_{x}\leq\pi$, where a prethermal plateau [25] is distinctly visible within the time steps $0\leq t/T\leq 100$. This is in sharp contrast to the behavior in one dimension, where magnetisation oscillations quickly decay, irrespective of the parameters $(\theta_{x},\theta_{z})$ away from the trivial point at $\theta_{x}=\pi$, as we have also confirmed in Supplementary Information S4 using the same quantum device.
We emphasise that the clean DTCs on the heavy-hexagonal lattice are quite stable in strongly interacting region $\theta_{J}=0.5\pi$ (see Supplementary Information S5). In Supplementary Information S7, we confirm that quantum scrambling [47, 48] has not yet occurred for $\theta_{x}\leq 0.8\pi$ at least by $t/T<50$. Since quantum scrambling is an underlying mechanism for the thermalisation of an isolated system [49, 50], the observed DTCs are stabilised in a prethermal regime. Compared to a square lattice with the same parameters, the speed of scrambling is much slower in a heavy-hexagonal lattice (see Supplementary Information S7). Furthermore, we indeed observe prethermal plateau structures in the magnetisation for small systems (see Supplementary Information S8), and thus conclude the emergence of prethermal DTCs for a parameter range of $0.8\pi\lesssim\theta_{x}<\pi$.